OpenStudy (anonymous):

what is the equation of the given line in standard form? the line is (-2, 0) (0, 4) (2, 8) A: 2x + y = 4 B: y = 2x +4 C: y - 4 =2x D: 2x - y = -4

OpenStudy (solomonzelman):

you rather mean that the line (the equation of which you are looking for) is going through those points: a) \(\large\color{black}{ \displaystyle \color{blue}{ (-2,0)} }\) b) \(\large\color{black}{ \displaystyle \color{green}{ (0,4)} }\) c) \(\large\color{black}{ \displaystyle \color{magenta}{ (2,8)} }\)

OpenStudy (solomonzelman):

( Usually, these questions only give you 2 points, but I think they are giving you 3 points, to make sure that there is no way to say that it is a non-linear function. )

OpenStudy (solomonzelman):

Anyway... do you know how to find the slope ?

OpenStudy (solomonzelman):

hey, you can reply... don't hesitate :)